OpenStudy (anonymous):

Suppose a tank has enough air to last 60 min at the surface. How long will it last at a depth of 99ft? (The pressure is 4 atm, or 4 times as great.

OpenStudy (anonymous):

PV = nRT P is pressure, n mols, R gas constant, T temperature, v Volume V = nRT/P Vsurface = nRT/P1 T = (Vsurface x P1)/nR) Vdepth = nRT/P2 T = (Vdepth x P2)/nr) Vsurface x P1 = Vdepth x P2 Vdepth = (Vsurface x P1)/P2 Vdepth = (A x 60 x 1)/4 (presuming volume is proportional to how long the air lasts with proportionality constant A) Vdepth = A60/4 = A15 approximately 15 minutes, depending on the constant A that links volume to how long it lasts

OpenStudy (anonymous):

How would you make a table showing how long the air will last at 0ft,20ft,33ft,40ft,50ft,66ft, and 99ft?

OpenStudy (anonymous):

nope, you can work it out. you need to work out the pressure at each depth, and then you can use the equation from before. at 99ft its 4atmos. so each ft adds 4/99 = 0.0404 atmospheres of pressure.

OpenStudy (anonymous):

How do I find the pressure on each different one tho?

OpenStudy (anonymous):

at 99ft pressure is 4 atmos. at 0ft pressure is 1atmos pressure difference is 3atmos so 3atmos/99ft = pressure per foot you know the feet (0, 20, 33 etc) so you can times that by the pressure per foot to get the pressure at each depth